OpenStudy (anonymous):

Geometry help please Malcolm bought a square sheet of paper of side 16 inches to decorate a pencil box he made in craft class. The pencil box has a height of 7 inches, a width of 4 inches, and a length of 5 inches. Approximately how many square inches of paper will be remaining if he pastes paper on to the bottom base and the four side surfaces of the pencil box without any overlap?

OpenStudy (hoblos):

area of paper = 16^2 =256 in^2 area to be covered = 2(7*4 + 7*5) + 4*5 = 146 in^2 remaining part = 256-146 = 110 in^2

OpenStudy (mertsj):

Bottom: 2(4 X 5)=40 front and back: 2(4 X 7)=56 right side and left side: 2(5 X 7)=70 Total: 40+56+70=146 Total sheet = 16X16 = 256 Left over: 256-146=110

OpenStudy (mertsj):